the width of a rectangle is 3 feet less than its length. if the area of the rectangle is 70 ft2, what are the dimensions of the rectangle.
Answer by Cromlix(4381)

Hi there,
Make the length = 'x'
width = x - 3
Area = Length x Width
70 = x(x - 3)
70 = x^2 - 3x
Sort into equation ax^2 + bx + c = 0
x^2 - 3x - 70 = 0
Factorise
(x + 7)(x - 10)= 0
x + 7 = 0
x = -7 (No answer as -ve)
.........
x - 10 = 0
x = 10
Length = 10ft
Width = 7ft.
Hope this helps:-)